# Characterization of FBK NUV-HD-Cryo SiPMs near LHe temperature

### v3

@@ IV Understanding the SiPMs’ temperature
In parallel to testing the SiPMs’ performance at 10 K, we launched a COMSOL [53] simulation to crosscheck whether the 10 K temperature measured by the sensor makes sense. In the simulation, all the materials and specifications are precisely the same as in the experimental setup; other input parameters are either directly measured or extrapolated from the measured data in situ. The simulation results are shown in Fig. 16. The temperature sensor near the SiPMs was 6.75 K, as the “T Sensor A (K)” indicates; while the SiPMs’ temperature was 7.1 K. That being said, there was a 0.35 K discrepancy between the temperature sensor and the SiPMs even though they have been mounted on the same copper plate and only have 5 cm distance. According to our experience in simulating other setups, the temperature discrepancy should be 0.1 K, given the two parts’ configuration. The “abnormal” discrepancy occurred because the SiPMs are connected to a 70 cm SMA cable, through which a preamplifier connected on another end at RT. As a result, the SiPMs has an extra “warm” source from the cable, in addition to the cryogenic source from the integrating sphere. Even by eye, we can see in Fig. 16 that the temperature on the cable’s cold end (near the SiPMs) is higher than the integrating sphere. In the figure, “PD” refers to the photodiode. “T Sensor B” is the temperature sensor near the PD, and “T Sensor C” is the temperature sensor on the cooling plate, which is the cryogenic source of the detector system.
@@ V Characterizing the SiPMs at 7 K
We re-measured the AP, CT, and PDE at 7 K again the same way as at 10 K. The combined results of AP at 7 K and 10 K are shown in Fig. 18. The plot has the following features: (a) the higher the bias voltage, the greater the AP; (b) the AP at 10 K is slightly higher ( 2 %) than 7 K (for an OV greater than 6 V); (c) for the SiPMs’ possible working bias voltage, OV = 10 V, the measured AP is 16% at 7 K and 10 K; and (d) with the 6 V OV, the AP we measured are 9% and 10.5% at 7 K and 10 K, respectively, which are consistent with reference [27]. Although a complete understanding of the AP at cryogenic temperatures is still lacking, reference [27] pointed out two conflicting processes that might help interpret the data. At a lower temperature, the SiPMs’ quenching resistance would increase, which will (i) suppress the avalanche triggering probability in recharge due to a lower voltage applied on the single photon avalanche diode (SPAD) and (ii) have a longer time constant to enhance the chance of a carrier to be released to generate an AP event. However, no known mechanism exists to characterize the AP phenomena precisely at low temperatures. We are communicating with FBK company to better understand our AP results.
Regarding the PDE tests, Fig. 19 does not show apparent temperature dependence at 7 K and 10 K. In addition, the plot shows that 405 nm light tends to generate a greater PDE than 530 nm, which is consistent with the tests on the similar type of SiPMs [34] at RT.
The current CT data collected at 7 K was odd. We need more time and data to understand it.
@@ VI.1 Discussion
VI.1.1 The FBK SiPMs’ charge distribution near LHe temperature
As shown in Fig. 7, the typical charge distribution of the FBK NUV-HD-Cryo SiPMs does not show separated Phe peaks at 10 K. We observed similar histograms at room temperature also. However, reference [34, 24, 27] all measured distinct Phe peaks at 77 K or RT on the similar types of FBK SiPMs, as summarized in table 2. We do not think our electronics and DAQ should be responsible for the unseparated Phe peaks because (a) as shown in Fig. 6, the noise of the single Phe signal is not significant, and (b) we tested a 4 4 mm2 Hamamatsu S13370-3050CN SiPMs at RT with almost the same electronics and observed clear peaks. After exchanging a couple of emails with FBK company, we are still far from concluding on this topic. We will continue to discuss this issue with FBK to fully understand it, and we hope to have the desired charge distribution eventually.
Type | Temp | Peaks | Win-size | Cell-pitch | Gain | Reference |
(K) | (Y/N) | (mm2) | (m2) | (E6) | ||
NUV-HD | 300 | Y | 11 | 2525 | 1.5 (OV 5 V) | [34] |
NUV-HD | 300 | Y | 11 | 4040 | 3.5 (OV 5 V) | [34] |
NUV-HD-LF | 77 | Y | 44 | 2525 | 0.5 (OV 5 V) | [27] |
NUV-HD-LF-HRq | 300 | Y | 11.77.9 | 2525 | 1.4 | [24] |
NUV-HD-Cryo | 7 | N | 11.77.9 | 3030 | 1.3 (OV 11 V) | This work |
S13370-3050CN | 300 | Y | 44 | 5050 | 2.6 (OV 4 V) | This work |
VI.1.2 Employ the FBK SiPMs into LHe TPCs
Although the lowest temperature the FBK NUV-HD-Cryo SiPMs have been verified to be functional so far is 7 K, instead of the liquid helium temperature of 4.5 K, due to the limitation of the cryocooler, we should be reassured about whether the SiPMs are eligible to work on an LHe TPC or not. The reason is that even if 7 K is the lowest temperature the FBK SiPMs can work, they can still be mounted on the TPC’s top and bottom bases as photosensors provided the SiPMs and the LHe were isolated by a 2-cm thick acrylic. Fig 20 indicates that a 2-cm thick acrylic would keep the SiPMs’ temperature at 7.2 K while maintaining an 89.3% transparency with the same assumption made by DarkSide-20k [24], a 1.5 cm acrylic layer would have a 90% transmission for TPB-converted light.

### v5

@@ IV Understanding the SiPMs’ temperature
In parallel to testing the SiPMs’ performance at 10 K, we launched a COMSOL [comsolWebsite] simulation to crosscheck whether the 10 K temperature measured by the sensor makes sense. In the simulation, all the materials and specifications are precisely the same as in the experimental setup; other input parameters are either directly measured or extrapolated from the measured data in situ. The simulation results are shown in Fig. 14. The temperature sensor near the SiPMs was 6.75 K, as the “T Sensor A (K)” indicates; while the SiPMs’ temperature was 7.1 K. That being said, there was a 0.35 K discrepancy between the temperature sensor and the SiPMs even though they have been mounted on the same copper plate and only have 5 cm distance. According to our experience in simulating other setups, the temperature discrepancy should be 0.1 K, given the two parts’ configuration. The “abnormal” discrepancy occurred because the SiPMs are connected to a 70 cm SMA cable, through which a preamplifier connected on another end at RT. As a result, the SiPMs has an extra “warm” source from the cable, in addition to the cryogenic source from the integrating sphere. Even by eye, we can see in Fig. 14 that the temperature on the cable’s cold end (near the SiPMs) is higher than the integrating sphere. In the figure, “PD” refers to the photodiode. “T Sensor B” is the temperature sensor near the PD, and “T Sensor C” is the temperature sensor on the cooling plate, which is the cryogenic source of the detector system.
@@ V Characterizing the SiPMs at 7 K
We re-measured the AP, CT, and PDE at 7 K again the same way as at 10 K. The combined results of AP at 7 K and 10 K are shown in Fig. 16. The plot has the following features: (a) the higher the bias voltage, the greater the AP; (b) the AP at 10 K is slightly higher ( 2 %) than 7 K (for an OV greater than 6 V); (c) for the SiPMs’ possible working bias voltage, OV = 10 V, the measured AP is 16% at 7 K and 10 K; and (d) with the 6 V OV, the AP we measured are 9% and 10.5% at 7 K and 10 K, respectively, which are consistent with reference [Acerbi17]. Although a complete understanding of the AP at cryogenic temperatures is still lacking, reference [Acerbi17] pointed out two conflicting processes that might help interpret the data. At a lower temperature, the SiPMs’ quenching resistance would increase, which will (i) suppress the avalanche triggering probability in recharge due to a lower voltage applied on the single photon avalanche diode (SPAD) and (ii) have a longer time constant to enhance the chance of a carrier to be released to generate an AP event. However, no known mechanism exists to characterize the AP phenomena precisely at low temperatures. We are communicating with FBK people to better understand our AP results.
Regarding the PDE tests, Fig. 17 does not show apparent temperature dependence at 7 K and 10 K. In addition, the plot shows that 405 nm light tends to generate a greater PDE than 530 nm, which is consistent with the tests on the similar type of SiPMs [Zappala16] at RT.
@@ VI.1 Discussion
Although the lowest temperature the FBK NUV-HD-Cryo SiPMs have been verified to be functional so far is 7 K, instead of the liquid helium temperature of 4.5 K, due to the limitation of the cryocooler, we should be reassured about whether the SiPMs are eligible to work on an LHe TPC or not. The reason is that even if 7 K is the lowest temperature the FBK SiPMs can work, they can still be mounted on the TPC’s top and bottom bases as photosensors provided the SiPMs and the LHe were isolated by a 2-cm thick acrylic. Fig 18 indicates that a 2-cm thick acrylic would keep the SiPMs’ temperature at 7.2 K while maintaining an 89.3% transparency with the same assumption made by DarkSide-20k [DarkSide20k17], a 1.5 cm acrylic layer would have a 90% transmission for TPB-converted light.